A hockey player scores a goal in 43% of his games. His team wins 65% of their games.In 31% of the team's games, the player scores a goal and the team wins. What is theprobability that the team wins if the player does not score a goal?
Answer #1

A - scores a goal

B - team wins their game

P(A) = 0.43

P(B) = 0.65

P(A and B) = 0.31

P(B| A') = P(A'and B) /P(A')

= (P(B) - P(A and B))/(1- P(A))

= (0.65 - 0.31)/(1- 0.43)

= 0.596491

Please rate